Suppose that 23 g of each of the following substances is initially at 26.0 ∘C. What is the final temperature of each substance upon absorbing 2.30 kJ of heat?

a.) gold t=________

b.) silver t =_______

c) aluminum t=_______

d) water _________
Answer #1

We know that Q= mC∆T or Q= mC(T2-T1), where, Q= heat absorbed , m= mass , C= specific heat, T1= Initial temperature and T2= Final temperature.

a. Gold

m= 23g, T1= 26°C ,Q= 2.30KJ =2300J and C= 0.129J/g °C

T2= (Q+CmT1)/Cm =( 2300J +0.129J/g °c * 23g* 26°c)/0.129J/g °c*23g = 801 °c

b. Silver

C= 0.240J/g °c

T2= (2300J+ 0.240J/g °c*23g* 26°c)/0.240J/g °c*23g= 443°c

C. Aluminium

C= 0.902J/g °c

T2= (2300+0.902*23*26)/0.902*23 = 137°c

D. Water

C= 4.179J/g°c

T2= (2300+4.179*23*26)/4.179*23= 50°c